OpenStudy (lucaz):

What are the steps for the derivative of 2^(3x)?

OpenStudy (dumbcow):

chain rule u =3x du = 3 \[\frac{d}{du} 2^u = \ln(2) 2^u\] so answer is: \[3 \ln(2) 2^{3x}\]

OpenStudy (lucaz):

I found this, but the answer is different..

OpenStudy (dumbcow):

they are equal just different forms 2^3 = 8

OpenStudy (lucaz):

like this (2^3)^x=2^(3x), just multiplying the exponents?

OpenStudy (dumbcow):

yes like that

OpenStudy (lucaz):

ok, thank you